OpenStudy (anonymous):

For the function f(x) = 2 – 3x, find f(–4). A. 14 B. 12 C. –10 D. –5

OpenStudy (anonymous):

@cwrw238 @texaschic101

OpenStudy (texaschic101):

f(x) = 2 - 3x f(-4) = 2 - 3(-4) f(-4) = 2 + 12 f(-4) = 14

OpenStudy (anonymous):

thanks a lot @texaschic101

OpenStudy (texaschic101):

no problem :)